OpenStudy (anonymous):

x^2-8x+30=0 finding roots??

OpenStudy (anonymous):

USE QUADRATIC FORMULA

OpenStudy (anonymous):

\[\frac{ -b \pm \sqrt{b^2 - 4ac} }{ 2a }\]

OpenStudy (anonymous):

I am for some reason in the wrong math class and I have no idea how to use the quadratic formula haha

OpenStudy (anonymous):

a = x^2 = 1 b = -8x = -8 c = 30 PLUG THEM INTO THE FORMULA

OpenStudy (anonymous):

so... a = 1 b = -8 c = 30

OpenStudy (anonymous):

here:

OpenStudy (anonymous):